Shift report 20010816 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 16 2001 - 00:00:10 EDT

  • Next message: Apache: "Shift report 20010816 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 16 00:00:09 2001, Shift: 16:00-24:00
    Supervisor: sjs, with: sjs, ah
    Log book#7 pages 96 - 97
    
    Shift Summary:
    ------------------------------------------------------------
    Beam cogged 16:00-17:30, 20:30-11:25
    Data runs 4220 (165K trigger1), 4223(198K trigger1)
    run 4220 has downscale of 3 for trigger 4
    run 4223 was not downscaled to allow for a total
    multiplicity measurement.
    
    FV adjusted scaler 1 gate to correct for a double pulsing
    situation, thereby reducing observed rate by about 1/3.
    
    RD came in at end of run 4223 and noted that tpm1 anode
    voltage was set too high (at 1230 V).  Stopped run to
    lower voltage to 1180 V.  Beam lost almost immediately
    after the run was stopped.
    
    brntuple not working at end of run 4220.  HI found problem 
    was associated with switch to root 3.01.06--he was able to
    fix the problem by the end of run 4223.  
    
    No new detector problems noted.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 16 2001 - 00:00:48 EDT